Fermi National Laboratory


Accelerator Update for August 17, 2005 - August 19, 2005

Wednesday, August 17
The day shift began with Operations monitoring store 4336 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter and MTest experiments taking beam.

Operations began preparing the Recycler for shot setup at 9:45 AM to stash antiprotons. Experts will use these shots to investigate the recent lost transfers.

The Recycler stash ended at 11:39 AM. None of the transfers were lost.

Pbar resumed stacking at 11:51 AM.

MCenter and MTest resumed taking beam at 12:02 PM.

At 2:33 PM, a Tevatron (TeV) RF station (TRF7) suffered a brief glitch, which caused the beam to grow into the abort gap; it recovered on its own. A TeV expert helped operators investigate.

A TeV RF expert turned off the heaters for TRF7 at 5:10 PM.

At 5:56 PM, a Booster RF station (BRF11) tripped off and caused the West Anode Power Supply (WAPS) to trip off also. Beam resumed at 6:02 PM, but then tripped when BRF11 tripped and caused WAPS to trip again. Operators bypassed BRF11 and called in an expert. He had the station back on by 7:30 PM. Beam resumed at 7:31 PM.

MTest experimenters reported at 10:55 PM that they were done with the beam for the night.

Operations terminated store 4336 at 11:02 PM.

TeV RF experts began replacing the modulator for TRF7 at 11:03 PM. They had the modulator changed about forty minutes later.

Operations put the TeV into shot setup at 11:50 PM.

Thursday, August 18
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established combination store 4337 at 1:49 AM with an initial luminosity of 106.91E30. Pbar resumed stacking and MCenter resumed taking beam.

Operations put the Recycler into shot setup at 5:03 AM. The stash ended at 6:09 AM. Pbar resumed stacking and MCenter resumed taking beam. Operations reported losing one transfer of antiprotons due to a Lambertson (LAM52) problem.

Recycler experts began a damper study at 10:20 AM.

At 11:29 AM, Operations reported that a lightning strike caused store 4337 to abort. It also caused a lot of other trips in the TeV, which were all eventually corrected.

Operation established beam to MCenter around noon.

Cryo system technicians began work at 12:06 PM on a wet engine in TeV sector D2. Then they moved to sectors B4, F3, and finally D3 for wet engine work. All the work was completed within two hours.

Operations put the TeV through a dry squeeze at 3:57 PM.

At 4:15 PM, Operations reported that a LCW leak was found in Meson coming from a magnet (MC6H1). This magnet was valved out. An expert determined that MCenter could run without this magnet.

Operations reported at 5:48 PM that they had put the TeV through a wet squeeze. Operations put the TeV through another wet squeeze at 6:43 PM.

Operations put the TeV into shot setup at 6:55 PM.

At 7:07 PM, Operations reported a Booster East Anode Power Supply (EAPS) tripped and wouldn’t reset. Operators investigated and discovered that BRF4 was the cause of the trip. Operators disconnected BRF4 from EAPS.

Operations established a combination store 4341 at 9:35 PM with an initial luminosity of 61.71E30. Operations reported that the poor luminosity was due to large transverse emittances and that was due to a kicker (A:EKIK) timing problem. Pbar resumed stacking.

MCenter resumed taking beam at 10:11 PM.

Friday, August 19
The midnight shift began with Operations monitoring store 4341, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing to stash antiprotons around 1:58 AM. Recycler finished stashing at 2:63 AM. Pbar resumed stacking.

MCenter resumed taking beam at 3:03 AM.

The Recycler stashed antiprotons at 6:45 AM and finished at 7:29 AM. Pbar resumed stacking and MCenter resumed taking beam.

The Plans for Friday and the Weekend
The plans for today and the weekend are to store, stack, and stash.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/19/2005   email Fermilab